# HiSAC: High-Resolution Sensing with Multiband Communication Signals

### v1

@@ 4.4. Temporal aggregation
Consider a sequence of $N$ path delays, $\{\hat{\tau}_{1}(t_{n}),\dots,\hat{\tau}_{L^{\rm OMP}}(t_{n})\}_{n=1}^{N}$, and amplitudes, $\{\hat{\alpha}_{1}(t_{n}),\dots,\hat{\alpha}_{L^{\rm OMP}}(t_{n})\}_{n=1}^{N}$, obtained by applying HiSAC to different ISAC packets or OFDM slots. These must be obtained in a short processing interval such that the underlying channel parameters can be considered constant. Recall that the delays outputted by OMP belong to a discrete grid of candidates $\nu_{q}$, with $q=1,\dots,Q$, which is kept fixed during the aggregation period. The temporal aggregation step is based on the following observation: When applied over a short time interval where the channel is almost constant, HiSAC outputs correlated sets of delays that can be aggregated (coherently or incoherently). To do so, Alg. 1 proceeds by first iterating over the elements in the delay grid and over time slots (line 2). Then, if $\nu_{q}$ is among the set of outputs of OMP in the considered slot $n$, we aggregate it to a running average of the path amplitudes, $\chi_{q}$ (lines 3-5). We propose two alternative versions of the algorithm for static targets and human sensing (line 4), respectively. If the targets are static, the temporal aggregation can be performed by taking into account the phase of the complex amplitudes (coherent aggregation), which gives higher Signal-to-Noise Ratio (SNR) and resolution. With dynamic targets such as humans instead, we only aggregate the magnitude information for each path (incoherent aggregation) since the time variation of the phase due to respiration or slight movement would lead to destructive interference. Finally, in line 13, we select as the final improved set of delays the $L^{\rm OMP}$ candidates for which $\chi_{q}$ is highest, which we denote by $\bar{\tau}_{l}$, for $l=1,\dots,L^{\rm OMP}$. The channel gains of such delays are the corresponding $\chi_{q}$, which we call $\bar{\alpha}_{l}$. The delays are then mapped to relative distances as $\bar{r}_{l}=c(\bar{\tau}_{l}-\bar{\tau}_{1})$.
@@
0:  Improved set of delays and amplitudes.
@@
2:  for $q=1,\dots,Q$ and $n=1,\dots,N$ do
3:     if $\nu_{q}\in\{\hat{\tau}_{l}(t_{n})\}_{l=1}^{L^{\rm OMP}}$ then
4:        $\chi_{q}\leftarrow\left\{\begin{array}[]{cc}\left[\hat{\alpha}_{l}(t_{n})+n\chi_{q}\right]/(n+1)&\mbox{ if target is static},\\ \left[|\hat{\alpha}_{l}(t_{n})|+n\chi_{q}\right]/(n+1)&\mbox{otherwise}.\end{array}\right.$
5:     end if
6:  end for
7:  Keep the $L^{\rm OMP}$ path delays with the highest $\chi_{q}$, with delays $\nu_{q}$.
@@ 5. Implementation
To implement HiSAC, we use the open-source Mimorph platform (Lacruz et al., 2021) as a baseline. The platform includes an AMD (Xilinx) RFSoC that comprises Field Programmable Gate Array (FPGA) logic, multiple AD/DA converters, and high-performance ARM processors. We implement HiSAC to work in the unlicensed 58-64 GHz mmWave band using linear antenna array front-ends with 16 elements from Sivers Semiconductors, suitable for analog beamforming. We choose a mmWave frequency band as it represents a challenging test case, given the high sensitivity to CFO and the strong phase noise of high frequencies (Rasekh et al., 2019). Signal conditioning from the RFSoC to the mmWave front-ends includes DC-block filters, low-pass filters (1 GHz cut-off frequency), and 3 dB attenuators. The main components of a TX-RX node of the testbed are shown in Fig. 5(a). We configure the testbed to work as $C=2$ incoherent subsystems with carrier frequencies 60.48 GHz and 62.64 GHz. The testbed can operate concurrently as TX and RX, in mono-static configuration, or as a bi-static system, using two nodes like the one in Fig. 5(a).
Signal generation. Signals for each sub-system are generated offline. Evaluating HiSAC requires collecting full band CFR estimates and other information used as ground truth, as detailed in Section 6.1. Therefore, we generate a composite packet including 5G-NR OFDM symbols and IEEE 802.11ay channel estimation fields. An OFDM symbol including De-Modulation-Reference Signal (DM-RS), spanning the full bandwidth, is used as ground truth (see the Full band baseline in Section 6.1). An IEEE 802.11ay Channel Estimation Field (CEF) is used for SC CIR estimation. 5G-NR pilot signals spanning different subbands are used for HiSAC, with bandwidth and starting frequency depending on the specific experiment. The signal is sent to the RFSoC using an Ethernet port and stored in loopback memories implemented in the FPGA logic. The FPGA clock is set to $245.76$ MHz with a super-sampling rate factor of 8, giving an equivalent of $1966.08$ MHz. The Inter-Frame Spacing (IFS) is configurable in runtime by a host PC. Since more than one subsystem is employed, independent data paths are used and connected to independent mmWave front-ends (Fig. 5(a)).
@@ 6.1. Experiments and baselines
Experiments description. To evaluate HiSAC, we perform 27 experiments involving ISAC radio channel measurements in different scenarios and configurations. Each experiment is repeated 5 times and involves the transmission of 100 packets with an IFS of 50 ms (unless stated otherwise). The experiments are carried out indoors, in a 7 m  $\times$ 6 m room (see Fig. 5(b)), and can be divided into six groups detailed in the following. In groups (1)-(5) we use metal cylindric reflectors as targets, while group (6) involves human subjects. Groups (1)-(4) and (6) are obtained in a mono-static scenario, while in (5) we use a bi-static setting. As anchor path to achieve phase coherence we use the self-interference path in the mono-static scenario and the LOS in the bi-static one.
@@
(4) Changing angle (5 experiments): 2 metal cylinders are placed about 2.5 m from the TX with 33 cm inter-target distance. We change the angular location of the targets in different experiments among $-30^{\circ},-15^{\circ},17^{\circ},30^{\circ}$. In each, experiment, we change the antenna Beam Pattern (BP) used by the TX to point at the targets. This scenario is of high practical interest since pilot signals transmitted in ISAC systems (e.g. SSBs) are often beamformed in different directions.
@@
(6) Human localization and tracking: A final group of experiments is performed with human targets, to demonstrate the effectiveness of HiSAC on weaker multipath components with respect to metal reflectors and its robustness to movement. These involve (i) 2 static subjects standing at 2.30 and 2.64 m from the TX as shown in Fig. 6(b), (ii) 2 moving targets walking back and forth from 3 to 1 m from the TX (in this case, the IFS is reduced to 5 ms). Experiments involving people have been carried out in compliance with the IRB of our institute and do not disclose information about the subjects.
Baselines for comparison. Since HiSAC is the first method to perform multiband ISAC, we evaluate it against the baseline methods described in the following. Note that (Wang et al., 2024), which is the closest prior work, is not suitable for comparison since (i) it is based on a deep neural network trained on sub-6 GHz signals, so it would need extensive data collection and retraining on our implementation, (ii) it uses a channel hopping scheme to collect CFR samples that is specific to sub-6 GHz Wi-Fi and does not apply to 5G-NR.
@@
SpotFi-based HiSAC (SpotFi-HiSAC). To demonstrate the effectiveness of HiSAC’s algorithm to achieve phase coherence across subbands, we design a competitor algorithm that uses SpotFi’s approach, (Kotaru et al., 2015), to perform this task. The rest of the delay estimation process is the same as in HiSAC, since the original SpotFi uses MUSIC to estimate delays which does not apply to non-contiguous subbands. SpotFi’s approach uses line fitting on the unwrapped phase of the CFR, then matches such models across the subbands, compensating for the differences in the slopes and intercepts.
@@ 6.2. In-depth evaluation
Carrier Aggregation. We start our evaluation with the CA scenario, combining multiple 5G-NR channels with $100$ and $400$ MHz bandwidth. We assume that in this case the CFR is estimated over the full channels, similar to using DM-RS pilots that span the full bandwidth (TS38.300, 2018). We consider two configurations with 2 subsystems:
@@
(i) Configuration 1 (CA-C1), including 4 subbands with an effective bandwidth of 400 MHz and a virtual bandwidth of 2.01 GHz. The first two subbands belong to subsystem 1, while the second two to subsystem 2. The starting frequencies of the subbands relative to the first one are $\{0,0.19,1.63,\\ 1.91\}$ GHz, while their bandwidths are all equal to 100 MHz;
@@
(ii) Configuration 2 (CA-C2), including 5 subbands with an effective bandwidth of 800 MHz and a virtual bandwidth of 3.46 GHz. The first three subbands belong to subsystem 1, while the second two to subsystem 2. The starting frequencies of the subbands relative to the first one are $\{0,0.19,1.2,2.88,3.36\}$ GHz, while their bandwidths are $\{0.1,\\ 0.1,0.4,0.1,0.1\}$ GHz.
@@
The main challenge in the CA scenario is to effectively combine subbands that are widely separated in the spectrum. Fig. 7 shows the RMSE and FRT obtained by HiSAC in experiment groups (1)-(3) with the CA use case. The horizontal dashed line represents the average RMSE obtained using the full band CFR. With CA-C1, HiSAC achieves accurate ranging with an average RMSE below 15 cm in all the experiments. The case with 3 targets gives the highest average RMSE due to the higher complexity of the multipath environment. The contiguous bandwidth and SpotFi-HiSAC yield comparable or worse RMSE. This proves that HiSAC gains ranging accuracy thanks to the increased virtual bandwidth and outperforms SpotFi’s method to achieve phase coherence. In terms of FRT, HiSAC provides significant gains. In the resolution limit test, the contiguous band is unable to resolve the targets (0.5 FRT), while HiSAC gives 0.93 FRT. With CA-C2, the overall performance in terms of RMSE and FRT on 2 and 3 targets improves for all methods due to the wider virtual bandwidth. However, only HiSAC significantly benefits from such increased virtual bandwidth when considering the resolution limit test, achieving higher FRT compared to CA-C1, while other methods perform slightly worse.
@@
Bandwidth Part. Next, we analyze a BWP scenario, where we combine multiple CFR estimates obtained by a 5G-NR system using SSB signals, spanning 240 subcarriers in the middle of the operating channel. Note that here the bandwidth of each subband is significantly lower than in the CA case since the CFR is measured only on a fraction of the total communication channel. Specifically, using 240 KHz subcarrier spacing, which is reasonable at mmWave frequencies, the bandwidth of each SSB is $B_{i,s}=57.6$ MHz. We consider two different configurations with 2 subsystems:
@@
(i) Configuration 1 (BWP-C1), including 4 subbands with an effective bandwidth of 460.8 MHz and a virtual bandwidth of 1.267 GHz. The first four subbands belong to subsystem 1, while the second four to subsystem 2. The starting frequencies of the subbands relative to the start of the full band are $\{0.02,0.12,0.22,0.32,0.91,1.01,1.11,1.21\}$ GHz;
@@
We stress that, as required by BWP operation, each CFR estimate in different subbands is obtained at a different time instant, with an IFS of 50 ms. Hence, the spectrum occupied at each time instant is just $57.6$ MHz. An example result on BWP-C1 is shown in Fig. 9, reporting the CFR and CIR squared magnitude ( Power-Delay Profile (PDP)). The main challenge in the BWP scenario is to combine narrow available subbands due to the use of unmodified SSB pilot signals. Fig. 8 shows the RMSE and FRT obtained by HiSAC in the experiment groups (1)-(3) with the BWP use case. With BWP-C1, HiSAC outperforms the other approaches. Note that with 3 targets it achieves slightly higher mean RMSE with respect to the contiguous band (2 cm), but has a higher FRT by 0.1. BWP-C2 represents a very challenging scenario due to the sparsity of the available spectrum (amounting to 18% of the virtual bandwidth) and to the narrow subbands used. In the resolution limit test, the contiguous CFR method fails in resolving the targets and has large RMSE, while HiSAC achieves 11 cm average RMSE and 0.67 FRT, with a resolution gain of 13% over SpotFi-HiSAC.
Figure 9. Example HiSAC results in BWP-C2. The top plot shows the subbands after coherency has been achieved and the reconstructed HiSAC CFR. The bottom plot shows the PDP.
@@
Impact of angle and beamforming. ISAC systems apply beamforming to direct the signal towards the communication RX or targets. HiSAC is robust to such changes in the direction of the transmission, as demonstrated by our results in Fig. 11, obtained on experiments from group (4). The target ranging RMSE remains below 15 cm when changing the angle from -30 to 30. The FRT does not change significantly when changing the transmission angle, proving that HiSAC can accurately measure distances in different directions.
Figure 11. HiSAC results changing target angle and transmission BP.
Bi-static setting. We evaluate HiSAC’s capability to estimate the targets’ distances in a bi-static setting, using experiments from group (5) and the BWP-C1/C2 configurations. Note that the full band range resolution in this case is reduced to $\Delta r=c/(2B\cos(\pi/4))\approx 17.5$ cm due to the bi-static angle being 90. Fig. 12 shows the RMSE and FRT obtained by HiSAC in the bi-static setting. These results are comparable to those obtained in the mono-static setting with a similar configuration, proving that HiSAC works when the TX and RX are widely separated. Conversely, the contiguous CFR can not resolve the targets with BWP-C2.
Figure 12. HiSAC results in a bi-static BWP scenario.
People localization. We test HiSAC on human subjects to demonstrate its capability to resolve weaker reflections, using the experiments from group (6). Fig. 13 shows the RMSE and the FRT. These results are obtained using the BWP-C1 and C2 configurations. HiSAC achieves 8 and 15 cm average RMSE with C1 and C2, respectively. Moreover, the FRT gain that it provides with respect to using the contiguous CFR is large: 0.25 using C1 and 0.17 using C2. Note that SpotFi-HiSAC’s FRT with human subjects degrades much more than HiSAC’s one. Being based on an anchor path, our method to achieve phase coherence is independent of the strength of the target multipath component.
Figure 13. HiSAC results with human subjects using BWP.
People tracking. Fig. 14 shows the PDP across time obtained from the CIR estimated by HiSAC and with the full band with 2 subjects walking back and forth. Peaks in the PDP (in yellow) correspond to the subjects. Although HiSAC’s PDP is noisier than that of the full band, its resolution, i.e., the capability of distinguishing the two subjects across time, is comparable. The bandwidth used by HiSAC in each time slot is only 57.6 MHz (20 times lower than the full bandwidth) corresponding to a resolution of 2.6 m.
Figure 14. Tracking 2 subjects across time in the BWP-C1 scenario. We plot the normalized PDP obtained from the reconstructed CIR with the full band (left) and HiSAC (right). HiSAC’s resolution is comparable to that of the full 1.2 GHz bandwidth.

### v2

@@ 4.4. Temporal aggregation
Consider a sequence of $N$ path delays, $\{\hat{\tau}_{1}(t_{n}),\dots,\hat{\tau}_{L^{\rm OMP}}(t_{n})\}$, and amplitudes, $\{\hat{\alpha}_{1}(t_{n}),\dots,\hat{\alpha}_{L^{\rm OMP}}(t_{n})\}$, for $n=1,\dots,N$, obtained by applying HiSAC to different ISAC packets or OFDM slots at time instants $t_{1},\dots,t_{N}$. These must be obtained in a short processing interval such that the channel parameters can be considered constant, i.e., $t_{N}-t_{1}$ should be within the coherence time of the channel.
Recall that the delays outputted by OMP belong to a discrete grid of candidates $\nu_{q}$, with $q=1,\dots,Q$, which is kept fixed during the aggregation period. The temporal aggregation step is based on the following observation: When applied within the coherence time of the channel, HiSAC outputs correlated sets of delays that can be aggregated (coherently or incoherently). To do so, Alg. 1 iterates over the elements in the delay grid and over time slots (lines 2-3). Then, if $\nu_{q}$ is among the set of outputs of OMP in the considered slot $n$, we accumulate it across time using a running average of the path amplitudes, $\chi_{q}$ (lines 4-6), that combines the current amplitude $\hat{\alpha}_{l}(t_{n})$ with the previous $\chi_{q}$.
We propose two alternative versions of the running average, one for static targets and one for human sensing (line 5), respectively:
  • If the targets are static, the temporal aggregation can be performed by taking into account the phase of the complex amplitudes (coherent aggregation), which gives higher Signal-to-Noise Ratio (SNR) and resolution. This corresponds to the first case in line 5 of Alg. 1.
  • With dynamic targets such as humans, instead, we only aggregate the magnitude information for each path (incoherent aggregation) since the time variation of the phase due to respiration or slight movement would lead to destructive interference. This corresponds to the second case in line 5.
Finally, in line 9, we select as the final improved set of delays the $L^{\rm OMP}$ candidates for which $\chi_{q}$ is highest, which we denote by $\bar{\tau}_{l}$, for $l=1,\dots,L^{\rm OMP}$. The channel gains of such delays are the corresponding $\chi_{q}$, which we call $\bar{\alpha}_{l}$. The delays are then mapped to relative distances as $\bar{r}_{l}=c(\bar{\tau}_{l}-\bar{\tau}_{1})$.
@@
0:  Improved set of delays and amplitudes $\{\bar{\alpha}_{l},\bar{\tau}_{l}\},l=1,\dots,L^{\rm OMP}$.
@@
2:  for $q=1,\dots,Q$ do
3:     for  $n=1,\dots,N$ do
4:        if $\nu_{q}\in\{\hat{\tau}_{l}(t_{n})\},l=1,\dots,L^{\rm OMP}$ then
5:           $\chi_{q}\leftarrow\left\{\begin{array}[]{cc}\left[\hat{\alpha}_{l}(t_{n})+n\chi_{q}\right]/(n+1)&\mbox{ if target is static},\\ \left[|\hat{\alpha}_{l}(t_{n})|+n\chi_{q}\right]/(n+1)&\mbox{otherwise}.\end{array}\right.$
6:        end if
7:     end for
8:  end for
9:  Return $\{\bar{\alpha}_{l},\bar{\tau}_{l}\},l=1,\dots,L^{\rm OMP}$ as the $L^{\rm OMP}$ path delays with the highest $\chi_{q}$.
@@ 5. Implementation
To implement HiSAC, we use the open-source Mimorph platform (Lacruz et al., 2021) as a baseline. The platform includes an AMD (Xilinx) RFSoC that comprises Field Programmable Gate Array (FPGA) logic, multiple analog-to-digital/digital-to-analog converters, and ARM processors. We implement HiSAC to work in the unlicensed 58-64 GHz mmWave band using linear antenna array front-ends with 16 elements from Sivers Semiconductors, suitable for analog beamforming. We choose a mmWave frequency band as it represents a challenging test case, given the high sensitivity to CFO and the strong phase noise of high frequencies (Rasekh et al., 2019). Signal conditioning from the RFSoC to the mmWave front-ends includes DC-block filters, low-pass filters (1 GHz cut-off frequency), and 3 dB attenuators. The main components of a TX-RX node of the testbed are shown in Fig. 5(a). We configure the testbed to work as $C=2$ incoherent subsystems with carrier frequencies 60.48 GHz and 62.64 GHz. The testbed can operate concurrently as TX and RX, in mono-static configuration, or as a bi-static system, using two nodes like the one in Fig. 5(a).
Signal generation. Signals for each sub-system are generated offline. Evaluating HiSAC requires collecting full band CFR estimates and other information used as ground truth, as detailed in Section 6.1. Therefore, we generate a composite packet including 5G-NR OFDM symbols and IEEE 802.11ay channel estimation fields. An OFDM symbol including De-Modulation-Reference Signal (DM-RS), spanning the full bandwidth, is used as ground truth (see the Full band baseline in Section 6.1). An IEEE 802.11ay channel estimation field is used for SC CIR estimation. 5G-NR pilot signals spanning different subbands are used for HiSAC, with bandwidth and starting frequency depending on the specific experiment. The signal is sent to the RFSoC using an Ethernet port and stored in loopback memories implemented in the FPGA logic. The FPGA clock is set to $245.76$ MHz with a super-sampling rate factor of 8, giving an equivalent of $1966.08$ MHz. The inter-packet time is configurable in runtime by a host PC. Since more than one subsystem is employed, independent data paths are used and connected to independent mmWave front-ends (Fig. 5(a)).
@@ 6.1. Experiments and baselines
Experiments description. To evaluate HiSAC, we perform 27 experiments involving ISAC radio channel measurements in different scenarios and configurations. Each experiment is repeated 5 times and involves the transmission of 100 packets with an inter-packet time of 50 ms (unless stated otherwise). The experiments are carried out indoors, in a 7 m  $\times$ 6 m room (see Fig. 5(b)), and can be divided into six groups detailed in the following. In groups (1)-(5) we use metal cylindric reflectors as targets, while group (6) involves human subjects. Groups (1)-(4) and (6) are obtained in a mono-static scenario, while in (5) we use a bi-static setting. In our experiments, we point the TX beam pattern toward the targets, unless stated otherwise. In practice, algorithms such as the one in (Pegoraro et al., 2024a, b) can be used to estimate the angular location of the scatterers. As an anchor path to achieve phase coherence, we use the self-interference path in the mono-static scenario and the LOS in the bi-static one.
@@
(4) Changing angle (5 experiments): 2 metal cylinders are placed about 2.5 m from the TX with 33 cm inter-target distance. We change the angular location of the targets in different experiments among $-30^{\circ},-15^{\circ},17^{\circ},30^{\circ}$. In each, experiment, we change the antenna beam pattern used by the TX to point at the targets. This scenario is of high practical interest since pilot signals in ISAC systems (e.g. SSBs) are often beamformed in different directions.
@@
(6) Human localization and tracking: A final group of experiments is performed with human targets, to demonstrate the effectiveness of HiSAC on weaker multipath components with respect to metal reflectors and its robustness to movement. These involve (i) 2 static subjects standing at 2.30 and 2.64 m from the TX as shown in Fig. 6(b), (ii) 2 moving targets walking back and forth from 3 to 1 m from the TX (in this case, the inter-packet time is reduced to 5 ms). Experiments involving people have been carried out in compliance with the IRB of our institute and do not disclose information about the subjects.
Baselines for comparison. Since HiSAC is the first method to perform multiband ISAC, we design the baseline methods described in the following. Note that (Li et al., 2024), which is the closest prior work, is not suitable for comparison since (i) it is based on a deep neural network trained on sub-6 GHz signals so that it would need extensive data collection and retraining on our implementation, (ii) it uses a channel hopping scheme to collect CFR samples that is specific to sub-6 GHz Wi-Fi and does not apply to 5G-NR.
@@
SpotFi-based HiSAC (SpotFi-HiSAC). To demonstrate the effectiveness of HiSAC’s algorithm to achieve phase coherence across subbands, we design a competitor algorithm that uses SpotFi’s approach, (Kotaru et al., 2015), to perform this task. SpotFi first extracts the phase of the CFR in each subband, representing it as a function of the subcarriers, and applies phase unwrapping to remove the wrapping around $2\pi$. Due to the linearity of the phase as a function of path delays, as can be seen from Eq. (1), a linear model can be fit to the unwrapped CFR of each subband. Then, the phase offsets are removed by applying a linear transformation to the phase of the CFRs. The parameters of the transformation are obtained by matching the slopes and intercepts of the linear models across subbands.
To fairly compare the effectiveness of the phase offsets compensation algorithm, we keep the rest of the delay estimation process the same as in HiSAC. This step is necessary since the original SpotFi uses the MUSIC algorithm to estimate delays, it can not handle non-contiguous subbands. Hence, SpotFi-HiSAC is an improvement of the original SpotFi that can be applied to non-contiguous subbands. Specifically, SpotFi-HiSAC applies, in order: the intra-subsystem OMP-based reconstruction from Section 4.1, the SpotFi phase offsets removal step, based on line fitting, and the multiband fusion step from Section 4.3.
@@ 6.2. In-depth evaluation
Carrier Aggregation. We start our evaluation with the carrier aggregation scenario, combining multiple 5G-NR channels with $100$ and $400$ MHz bandwidth. We assume that in this case the CFR is estimated over the full channels, similar to using DM-RS pilots that span the full bandwidth (TS38.300, 2018). We consider two configurations with 2 subsystems:
@@
(i) Configuration 1 (CA-C1), including 4 subbands with an effective bandwidth of 400 MHz and a virtual bandwidth of 2.01 GHz. The first two subbands belong to subsystem 1, while the second two to subsystem 2. The starting frequencies of the subbands relative to the first one are $\{0,0.19,1.63,1.91\}$ GHz, while their bandwidths are all equal to 100 MHz;
@@
(ii) Configuration 2 (CA-C2), including 5 subbands with an effective bandwidth of 800 MHz and a virtual bandwidth of 3.46 GHz. The first three subbands belong to subsystem 1, while the second two to subsystem 2. The starting frequencies of the subbands relative to the first one are $\{0,0.19,1.2,2.88,3.36\}$ GHz, while their bandwidths are $\{0.1,0.1,0.4,0.1,0.1\}$ GHz.
@@
The main challenge in the carrier aggregation scenario is to effectively combine subbands that are widely separated in the spectrum. Fig. 7 shows the RMSE and FRT obtained by HiSAC in experiment groups (1)-(3) with the carrier aggregation use case. The horizontal dashed line represents the average RMSE obtained using the full band CFR. With CA-C1, HiSAC achieves accurate ranging with an average RMSE below 15 cm in all the experiments. The case with 3 targets gives the highest average RMSE due to the higher complexity of the multipath environment. The contiguous bandwidth and SpotFi-HiSAC yield comparable or worse RMSE. This proves that HiSAC gains ranging accuracy thanks to the increased virtual bandwidth and outperforms SpotFi’s method to achieve phase coherence. In terms of FRT, HiSAC provides significant gains. In the resolution limit test, the contiguous band is unable to resolve the targets (0.5 FRT), while HiSAC gives 0.93 FRT. With CA-C2, the overall performance in terms of RMSE and FRT on 2 and 3 targets improves for all methods due to the wider virtual bandwidth. However, only HiSAC significantly benefits from such increased virtual bandwidth when considering the resolution limit test, achieving higher FRT compared to CA-C1, while other methods perform slightly worse.
@@
Bandwidth Part. Next, we analyze a bandwidth part scenario, where we combine multiple CFR estimates obtained by a 5G-NR system using SSB signals, spanning 240 subcarriers in the middle of the operating channel. Note that here the bandwidth of each subband is significantly lower than in the carrier aggregation case since the CFR is measured only on a fraction of the total communication channel. Specifically, using 240 KHz subcarrier spacing, which is reasonable at mmWave frequencies, the bandwidth of each SSB is $B_{i,s}=57.6$ MHz. We consider two different configurations with 2 subsystems:
@@
(i) Configuration 1 (BWP-C1), including 4 subbands with an effective bandwidth of 460.8 MHz and a virtual bandwidth of 1.267 GHz. The first four subbands belong to subsystem 1, while the second four to subsystem 2. The starting frequencies of the subbands relative to $f_{0}$ are $\{0.02,0.12,0.22,0.32,0.91,1.01,1.11,1.21\}$ GHz;
@@
We stress that, as required by bandwidth part operation, each channel estimate in different subbands is obtained at a different time instant, with an inter-packet time of 50 ms. Hence, the spectrum occupied at each time instant is just $57.6$ MHz. An example result on BWP-C1 is shown in Fig. 9, reporting the CFR (above) and the squared magnitude of the CIR (below). The main challenge in the bandwidth part scenario is to combine narrow available subbands due to the use of unmodified SSB pilot signals. Fig. 8 shows the RMSE and FRT obtained by HiSAC in the experiment groups (1)-(3) with the bandwidth part use case. With BWP-C1, HiSAC outperforms the other approaches. Note that with 3 targets it achieves a slightly higher mean ranging error with respect to the contiguous band (2 cm), but resolves a higher fraction of targets by 0.1. BWP-C2 represents a very challenging scenario due to the sparsity of the available spectrum (amounting to 18% of the virtual bandwidth) and to the narrow subbands used. In the resolution limit test, the contiguous CFR method fails in resolving the targets and has a large ranging error, while HiSAC achieves 11 cm average error and resolves 67% of the targets on average, with a resolution gain of 13% over SpotFi-HiSAC.
Figure 9. Example HiSAC results in BWP-C2. The top plot shows the subbands after coherency has been achieved and the reconstructed HiSAC CFR. The bottom plot shows the CIR squared magnitude.
@@
Impact of angle and beamforming. ISAC systems apply beamforming to direct the signal towards the communication RX or targets. HiSAC is robust to such changes in the direction of the transmission, as demonstrated by our results in Fig. 11, obtained on experiments group (4). The ranging RMSE remains below 15 cm when changing the angle in $[-30^{\circ},30^{\circ}]$. The FRT does not change significantly when changing the transmission angle, proving that HiSAC accurately measures distances in different directions.
Figure 11. HiSAC results changing target angle and transmission beam pattern.
Bi-static setting. We evaluate HiSAC’s capability to estimate the targets’ distances in a bi-static setting, using experiments from group (5) and the BWP-C1/C2 configurations. Note that the full band range resolution in this case is reduced to $\Delta r=c/(2B\cos(\pi/4))\\ \approx 17.5$ cm due to the bi-static angle being 90. Fig. 12 shows the RMSE and FRT obtained by HiSAC in the bi-static setting. These results are comparable to those obtained in the mono-static setting with a similar configuration, proving that HiSAC works when the TX and RX are widely separated. Conversely, the contiguous CFR can not resolve the targets with BWP-C2.
Figure 12. HiSAC results in a bi-static bandwidth part scenario.
People localization. We test HiSAC on human subjects to demonstrate its capability to resolve weaker reflections, using the experiments from group (6). Fig. 13 shows the RMSE and the FRT. These results are obtained using the BWP-C1 and C2 configurations. HiSAC achieves 8 and 15 cm average RMSE with C1 and C2, respectively. Moreover, the FRT gain that it provides with respect to using the contiguous CFR is large: 0.25 using C1 and 0.17 using C2. Note that SpotFi-HiSAC’s FRT with human subjects degrades much more than HiSAC’s. Being based on an anchor path, our method to achieve phase coherence is independent of the strength of the target multipath component.
Figure 13. HiSAC results with human subjects using bandwidth part.
People tracking. Fig. 14 shows the CIR squared magnitude across time obtained from the CIR estimated by HiSAC and with the full band with 2 subjects walking back and forth. Peaks in the CIR (in yellow) correspond to the subjects. Although HiSAC’s CIR is noisier than that of the full band, its resolution, i.e., the capability of distinguishing the two subjects across time, is comparable. The bandwidth used by HiSAC in each time slot is only 57.6 MHz (20 times lower than the full bandwidth), i.e., 2.6 m resolution.
Figure 14. Tracking 2 subjects across time in the BWP-C1 scenario. We plot the normalized CIR squared magnitude obtained from the reconstructed CIR with the full band (left) and HiSAC (right). HiSAC’s resolution is comparable to that of the full 1.2 GHz bandwidth.